OpenStudy (anonymous):

Kalon and his friend Marna own a chimney sweep service company. Working together, they can clean a chimney in 1 5/7 hours. If it takes Kalon 4 hours to clean a 20-foot chimney by himself, how long does it take Marna to clean the same size chimney by herself?

OpenStudy (anonymous):

\[\frac{ kalon }{ h } = chimney\]\[\frac{ Marna}{ h } = chimney\]the hours they need are not the same, right?

OpenStudy (anonymous):

no i don't think so.

OpenStudy (anonymous):

we can write how fast they are in terms of 1 chimney. \[Kalon:\frac{ chimney }{ 4h }\]\[Marna:\frac{ chimney }{ ?h }\]

OpenStudy (anonymous):

if we want to add them together (Working together) we must get a common denominator. you will see this in the formula that is used for problems like this

OpenStudy (anonymous):

\[Kalon:\frac{ chimney \times ?h }{ 4h \times ?h}\]\[Marna:\frac{ chimney \times 4h}{ ?h \times 4h}\]

OpenStudy (anonymous):

do you have some idea where we get this from so far? :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

great

OpenStudy (anonymous):

kalon: 1/4 ( 1 chimney per 4 hours) marna: 1/x <- we don't know how long it takes for marna to make a chimney yet 1/4 + 1/x = the amount of chimneys they do together

OpenStudy (anonymous):

verify if this is correct

hero (hero):

The working together formula is \(\dfrac{K \times M}{K + M} = t\) K = amount of hours Kalon can clean the chimney alone M = amount of time Marna can clean the chimney alone. t = the time both can clean the chimney together.

OpenStudy (anonymous):

thanks Hero :)

OpenStudy (anonymous):

I knew there was this equation, I couldn't remember it from my mind

hero (hero):

In this case, \(t = 1 + \frac{5}{7}\) in hours \(M = 4\) in hours We need to find K

hero (hero):

For convenience, we can express \(t\) as \(\dfrac{12}{7}\)

OpenStudy (anonymous):

@AlexaLeyva I hope you can see better why the formula Hero uses works :) it is related to the work we have done

OpenStudy (anonymous):

yeah it is. thanks anyways @mathessentials

OpenStudy (anonymous):

thank you @Hero

hero (hero):

So to set up the equation we have: \(\dfrac{4K}{4 + K} = \dfrac{12}{7}\) \(28K = 12(4 + K)\) \(28K = 48 + 12K\) \(28K - 12K = 48\) \((28 - 12)K = 48\) \(16K = 48\) \(K = \dfrac{48}{16}\) \(K = 3\) So Karna can clean the chimney alone in 3 hours

OpenStudy (anonymous):

thats what i got